OpenStudy (anonymous):

Suppose that for some values of A and r, P(t)=Ae^rt and we know that P(8)=175 and P(11)=125. Find A and r.

OpenStudy (anonymous):

175 = A * e^8r 125 = A * e^11r Divide: => 125/ 125 = e^ -3r Take ln: => - 3r = ln125/ 175 => r = .112 = 11.2 % Thus A = 125/ e^11r = 125/3.43 = $36.4